a person travels a total of 35 minutes. They jog to a location at 5 meters per second and walk back at 2 meters per second. How many total meters did they travel?
Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Ok, lets deal in minutes:
5 meters/sec=5*60 =300 meters/min
2 meters/sec=2*60=120 meters/min
Let d=one-way distance
t1=time required jogging out
t2=time required walking back
t1+t2=35
t1=d/300
t2=d/120
d/300+d/120=35 multiply each term by 600
2d+5d=21000
7d=21000
d=3000 meters
2d=6000 meters total meters travelled
Hope this helps--ptaylor